OpenStudy (anonymous):

inverse of (3x-10)/(9x+9)

hero (hero):

y = (3x-10)/(9x+9) x = (3y-10)/(9y+9) x(9y+9) = 3y-10 9xy+9x = 3y - 10 9x - 10 = 3y - 9xy 9x - 10 = y(3 - 9x) 9x - 10/(3 - 9x) = y inverse of x